OpenStudy (anonymous):

Suppose you roll a six-sided die two times hoping to get two numbers whose sum is even. What is the sample space? How many favorable outcomes are there?

OpenStudy (anonymous):

wouldnt it be 18/36

OpenStudy (anonymous):

i just need clarification

OpenStudy (perl):

@AriPotta those possibilities are not equally likely

OpenStudy (perl):

it is true that the sample space is { 2,3,4,5,6,7,8,9,10,11,12} , but the sample points are not equally likely.

OpenStudy (perl):

you want to know probability of rolling a sum of 2,4,6,8,10,12 you can use a diagram of the sum of two die , we can google an image

OpenStudy (perl):

P( roll 2) = 1/36 P(roll 4) = 3/36 P( roll 6) = 5/36 P(roll 8) = 5/36 P( roll 10) = 3/36 P( roll 12) = 1/36 now add up these probabilities

OpenStudy (anonymous):

18/36?

OpenStudy (perl):

or 1/2

OpenStudy (anonymous):

would that be the sample space or number of favorable outcomes

OpenStudy (perl):

I used the modified sample space , weighting the sample points

OpenStudy (anonymous):

so what would be the number of favorable outcomes?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

if you want to use the original 36 point sample space, then you would have 18 favorable, and 36 total

OpenStudy (perl):

original 36 point sample space S = { (1,1) (1,2) (2,1) (1,3) (2,2) (3,1) ...}

OpenStudy (perl):

then the favorable events are (1,1) (1,3) (2,2) (3,1) ... those that add to an even number